OpenStudy (anonymous):

A plane trip took 6 hours at a speed of 360 mph. How far did the plane go?

OpenStudy (unklerhaukus):

\[6\text h\times 360 \text{mi/h}=6\times360 \text{mi}\]

Parth (parthkohli):

\[\LARGE => 360 * 6 = 2160 miles\]

OpenStudy (anonymous):

360mph x 6hours = 2160 miles

OpenStudy (radar):

Assuming that is ground speed. d=st d=distance s=speed t=time with all being in similar units like. distance in miles, speed in miles per hour, and time in hours.

OpenStudy (anonymous):

space is speed*time, so 6*360 = 2160. Just using dimensional analysis you could have the solution. In fact h * mi/h = h/h * mi = mi that is the physical quantity of space.

OpenStudy (anonymous):

I use the following setup for any problem like this (that has miles per hour, or a conversion of minutes or seconds or anything). Look at the file. All you have to do is add in your amounts given (per hour = / 1 hour) and do the simple math.

OpenStudy (anonymous):

thanks yall